# Word-specific tonal realizations in Mandarin

### v1

@@ 2.3.4 Segment-related controls
Onset: The effect of onset consonant on F0 has been studied in considerable detail in Mandarin. Ho, (1976), for example, found that voiced onsets lead to lower F0 in the following vowel, as compared to voiceless onsets. Aspiration, on the other hand, often causes pitch perturbation, resulting in higher F0, although the magnitude of this effect appears to be tone dependent (Xu and Xu, , 2003). Following Howie, (1974), we distinguished onset types according to manner of articulation, voicing, and aspiration. For each of the two syllables in our target words, we distinguished between ‘aspirated-affricate’, ‘aspirated-stop’, ‘unaspirated-affricate’, ‘unaspirated-stop’, ‘voiceless-fricative’, and ‘voiced’. Syllables that do not have onsets and start with vowels or glides instead were coded as ‘null’. Our dataset contained 30 different combinations of the onset type of the first syllable (onset1) and that of the second syllable (onset2).
@@ 2.4 Modeling Strategy
In GAMs, the issue of autocorrelation can be addressed by incorporating a first order autoregression model for the errors, denoted as AR(1). An AR(1) model is a linear model that predicts a given value of a time series from the immediately prior value; including an AR(1) process in a GAM enables the model to accommodate structure in the residuals by positing a linear relationship between a given residual and its preceding residual. This can be a highly effective way of dealing with autocorrelation. However, for the current dataset, including an AR(1) process led to two new problems (see Appendix A for further details). Firstly, the AR(1) process in a GAM assumes that the degree of autocorrelation is invariant; however, the degree of autocorrelation actually varies considerably across the tokens in our dataset and the AR(1) process therefore over-corrects for some tokens but under-corrects for others. Secondly, the AR(1) process has an undesirable effect on the overall distribution of the residuals in our models, increasing the extent to which they deviate from normality; this happens because the data contains a large number of discontinuous contours arising from voicelessness or creaky voice (see Figures A3 and A4). For these two reasons, we decided not to incorporate an AR(1) process in our modeling.
The decision not to include an AR(1) process means that we have fewer independent datapoints than a GAM assumes. Our models are therefore anti-conservative and calculate confidence intervals that are too narrow. For this reason, in the analyses that follow, we do not report p-values for the partial effects in our GAMs. Instead, we adopted a two-fold modeling strategy. Firstly, in order to evaluate whether a given predictor is relevant for understanding F0 contours, we made use of Akaike’s Information Criterion (AIC), assessing the extent to which AIC decreases, i.e. model fit improves, when a predictor is added to a baseline model. Secondly, we investigated the adequacy of our predictors by means of cross validation. That is, we held out a small portion of our dataset as testing data, fitted our models to the remaining data, and then assessed model accuracy on the testing data. In this way, we can assess the precision with which our models predict novel, previously unseen, data, and establish whether inclusion of a predictor improves prediction accuracy.
@@
    Variation in the tonal realization of a word cannot be reduced to the segment-related constraints on articulation previously described in the literature.
@@
    Information about a word’s meaning in context will improve prediction of its tonal realization.
@@
Prediction 1 is based on the hypothesis that the unique pitch contour of each spoken Mandarin word token is determined in part by the meaning of that token. If this is correct, then variations in tonal realization arise not only from the mechanical constraints on articulation previously described in the literature, but also as a result of form-meaning connections in the lexicon. We therefore investigated whether using word as a predictor would lead to a more precise model of tonal realization, as compared to a model using the set of segment-related predictors (e.g., vowel height) that have previously been identified as relevant. Since word incorporates not only the segmental form of a word type but also the associated semantics, our expectation was that it would be a superior predictor compared to all the previously identified segment-related variables considered jointly, when we controlled for contextual effects such as speech rate.
@@
As a first step, we fitted a baseline GAM to log-transformed F0, including all the speaker-related and context-related control variables described in Sections 2.3.2 and 2.3.3, in interaction with time. We then fitted six further models, each with one segment-related control variable added to the baseline, allowing us to investigate the articulatory effects described in the literature. To compare these effects jointly against the effect of word, we fitted two additional models, one with all six segment-related control variables included in addition to the baseline, and the other with only word as an additional predictor. Both word and the segment-related controls (vowel1, vowel2, onset1, onset2, syllable1, syllable2) were modeled as factor smooths in interaction with time.
@@
Prediction 2 is based on the hypothesis that variation in tonal realization serves to make word tokens with different meanings more discriminable from one another. To address this prediction, we compared the model with word as the sole predictor added to the baseline against a model in which sense was the sole predictor added to the baseline. Our expectation was that if semantics is indeed at issue, replacing word by sense should improve model fit even further. In the dataset for the word-type analysis, the frequency distribution of sense is skewed towards the right, with about half of the senses having no more than 13 tokens. To make sure that all senses included in the models had sufficient tokens for statistical evaluation, we therefore used only a subset of the data for the models used to investigate the effect of sense. Since the median number of tokens per sense in the dataset was 13.5, we excluded senses with fewer than 14 tokens. This left us with a dataset of 3,458 tokens representing 65 senses across a smaller set of 48 word types. We used this smaller dataset for models evaluating sense as a predictor. [^12] For completeness, we note that 35 words in the smaller dataset have only one sense. Of these 35 words, six have only one sense listed in the Chinese Wordnet, and seven are not included in the vocabulary of the Chinese WordNet. For the rest, either the tagger only identified one sense, or only one sense of the word had more than 13 tokens in our dataset. The statistical analysis proceeded in the same way as described above for word type, except that we added the additional model with sense as the only predictor in addition to the baseline.
We fitted our models using the mgcv package (Wood, , 2017) in R (R Core Team, , 2022). The distributions of the residuals of Gaussian models fitted to the pitch contours were t-distributed rather than normally distributed, and we therefore set the family argument to scat (scaled-t, Wood et al., , 2016).
@@ 2.5 The baseline GAM
We fitted the baseline GAM to log-transformed F0, including all speaker-related and context-related predictors, using the following model specification:
@@
s(speech_rate) + ti(time, speech_rate) +
@@
The first line of this model requests a main effect for gender, in order to account for male voices being lower on average than female voices. In addition, the model requests a separate smooth for each gender. The upper left-hand panel of Figure 3 plots the predicted contours for speakers identified as female (red) and those identified as male (blue). Similar to the pattern observed in read speech (cf. Figure 1), the realization of the RF tonal pattern in spontaneous speech is characterized by a shallow fall, followed by a long rise, and finally a much larger fall. Speakers identified as male show reduced pitch excursion compared to those identified as female, presumably due to male voices having a more compressed pitch range. The second line of the model requests by-speaker nonlinear random effects, using factor smooths. These factor smooths specify, for each speaker, the specific way in which that particular speaker modulates the general F0 contour associated with their gender. [^13] The syntax bs=’fs’ on the second line of the baseline model specification has a similar effect to the by argument on the first line; both terms request a smooth for each level of a single factor variable (in this case, speaker and gender respectively). The by argument is generally preferred when the factor has few levels, and the levels are of interest, as in the case of gender. For computational efficiency, the fs argument is preferred when there are many levels and these levels are of less direct interest, as in the case of speaker. When specifying a by-smooth, a separate term requesting a main effect for the intercept needs to be specified. In contrast, a factor smooth incorporates adjustments to the intercept, thus effectively calibrating the individual smooths for their relative position with respect to the general intercept.
@@
The next four lines in the model formula deal with the four numerical context-related controls, namely speech_rate, utterance_position, bigram_previous and bigram_following. For each of these variables, the model requests a main effect smooth in combination with a tensor product smooth for the interaction of the given variable with time (using an interaction-specific tensor product smooth specified with ti). The upper mid panel of Figure 3 plots the modulating effect of speech rate on the base contour of speakers identified as female. Higher speech rates, represented by lighter shades of red, reduce the amplitude of the wave. The effect of position in the utterance is depicted in the upper right-hand panel of Figure 3. The tonal shape is clearly most different when the word occurs towards the end of an utterance, in which case we observe an earlier peak. This might be due to the fact that words in singleton utterances are treated as occurring at the end of an utterance. The left and mid panels in the lower row of Figure 3 present the effects of the bigram probabilities given the preceding and following word respectively. Similar to the effect of speech rate, higher bigram probabilities, represented by lighter shades of red, generally lead to a smaller F0 excursion.
@@
Figure 3: Baseline GAM. The upper left-hand panel shows the predicted base contours for speakers identified as female (red) and speakers identified as male (blue). The next four panels show, for female speakers, how the base contour is modulated by speech rate, utterance position, previous bigram probability, and following bigram probability, respectively. The final panel presents, again for female speakers, the effect of tonal coarticulation witWoh the tone of the preceding word, given that the following word has a high-level tone.
@@ 2.6.1 Evaluation of predictors
The left-hand panel of Figure 4 presents the improvement in model fit, as compared to the baseline model, for the six models with a single additional segment-related control, the model with all segment-related controls (henceforth the omnibus-segment model), and the model with only word as additional predictor. Improvement in model fit is gauged by the magnitude of any decrease in AIC. As can be seen, all segment-related controls improve model fit, a result that dovetails well with previous studies.
@@
The inclusion of all six segment-related controls jointly in the omnibus-segment model provides a substantially better fit than any single predictor. However, adding only word to the baseline model results in an even better fit, the difference in AIC being no less than 5804 AIC units. Although the segment-related controls address all the word properties previously found to influence tonal realization, the contribution of just word by itself to the model fit is much stronger. It is clear that the association between word type and tonal realization cannot be reduced to the segment-related constraints on articulation previously described in the literature. The actual pitch contour is richer than what can be predicted from all phonetic features that have been found to be relevant.
The omnibus-segment model has the problem that its key predictors are correlated with one another. In a GAM, the nonlinear equivalent of collinearity is concurvity. The concurvity score of a predictor is a number bounded between zero and one that measures the degree to which the effect of a given independent variable can be predicted by one or more of the other independent variables in the model. If a predictor’s concurvity is low, this predictor has its own explanatory value. However, if the concurvity is high, the predictor is strongly confounded with other predictors.
The right-hand panel of Figure 4 presents the concurvity scores of the segment-related controls in the omnibus-segment model (marked in blue) and the concurvity score of the predictor word in the word-type model (red). For the omnibus-segment model, concurvity scores of all predictors are high, which is perhaps unsurprising given the restrictive phonotactic constraints governing Mandarin syllables (Duanmu, , 2007). The high concurvity indicates that the effects of the segment-related controls are confounded with one another, rendering interpretation of the individual effects difficult if not impossible. On the other hand, in the word-type model, the concurvity of word is low, so that the interpretation of the effect of individual word types on the F0 contours is straightforward.
@@
Figure 4: Word-type analysis. The left-hand panel shows model fit improvement, gauged by decrease in AIC units when a predictor (or set of predictors) is added to the baseline model. The right-hand panel shows the concurvity score of individual predictors in two models: the omnibus-segment model with all segment-related control variables (blue) and the word-type model with word as predictor (red).
@@
The predicted pitch contours of a sample of 15 word types are presented in Figure 5. To better visualize how the word-specific tonal modulations differ from one another, the partial effect predicted for each word has been added to the general contour for speakers identified as female (cf. Figure 3). In general, the fall-rise-fall pattern can be observed for all these words, but the details of tonal excursions differ significantly from word to word. For example, while the initial falling part is very prominent for words like (c) {CJK*}UTF8bsmi決定 juédìng ‘decision’ and (f) {CJK*}UTF8bsmi全部, quánbù ‘all’, it is rather muted for (m) {CJK*}UTF8bsmi一半 yíbàn ‘half’ and (d) {CJK*}UTF8bsmi年紀 niánjì ‘age’. In addition, in terms of the degree of undulation, some words have more reduced tonal range, such as (a) {CJK*}UTF8bsmi不是 búshì ‘not’ as compared to (j) {CJK*}UTF8bsmi文化 wénhuà ‘culture’, which has an extensive F0 excursion.
@@
Figure 5: Word-type analysis. Examples of the pitch contours predicted by the general smooth for time for female speakers, combined with the partial effects of the factor smooth for word. These partial effects do not include the general intercept, nor the differences in pitch between female and male speakers. As they represent the pure effect of word on the pitch contour, irrespective of other predictors, the curves are centered around the y-axis (indicated by a horizontal dotted line). The vertical dotted lines in the panels indicate the average (word-specific) syllable boundary.
@@
2.6.2 Cross validation
We have seen that word type is an excellent predictor of pitch realization, outperforming the segment-related predictors previously identified in the literature. However, due to the presence of strong temporal autocorrelations inherent in the pitch measurements, the confidence intervals in our models, shown for example in Figure 5, are anti-conservative and are therefore far from optimal for estimating the uncertainty of the predicted partial effects. Nevertheless, the models do make explicit predictions of F0 contours, and can do this even for held-out data, i.e. tokens that were withheld from the model during model fitting (training). If our hypothesis is correct, a GAM that has access to word should provide superior prediction accuracy on held-out data compared to a GAM that has access only to the segment-related controls. We therefore evaluated prediction accuracy under cross validation. We first held out 10% of the current data as test data, and used the remaining 90% as training data. We ensured that every word type was represented in both the training and test data, and that the number of tokens per type in the test data was proportional to that in the training data.

### v2

@@ 2.3.4 Segment-related controls
Onset: The effect of onset consonant on F0 has been studied in considerable detail in Mandarin. Ho, (1976), for example, found that after voiced consonants, pitch tends to start lower than after voiceless consonants. For stops, aspiration also results in lower initial F0, although the magnitude of this effect appears to be tone dependent (Xu and Xu,, 2003). Following Howie, (1974), we distinguished onset types according to manner of articulation, voicing, and aspiration. For each of the two syllables in our target words, we distinguished between ‘aspirated-affricate’, ‘aspirated-stop’, ‘unaspirated-affricate’, ‘unaspirated-stop’, ‘voiceless-fricative’, and ‘voiced’. Syllables that do not have onsets and start with vowels or glides instead were coded as ‘null’. Our dataset contained 30 different combinations of the onset type of the first syllable (onset1) and that of the second syllable (onset2).
@@ 2.4 Modeling Strategy
In GAMs, the issue of autocorrelation can be addressed by incorporating a first order autoregression model for the errors, denoted as AR(1). An AR(1) model is a linear model that predicts a given value of a time series from the immediately prior value; including an AR(1) process in a GAM enables the model to accommodate structure in the residuals by positing a linear relationship between a given residual and its preceding residual. Specifically, the residual at time $t$ is modeled as the sum of a proportion $\rho$ of the residual at time $t-1$ plus Gaussian noise. In all the models presented below, the AR(1) process is included. In order to determine the appropriate value of $\rho$, we first fitted a model without AR(1), then calculated the autocorrelation of the residuals in this model at lag 1, i.e., comparing the residual at each time point with the immediately preceding residual. Finally we fitted the model again now with AR(1), for which we set $\rho$ to the lag 1 autocorrelation. [^14] We note that the current implementation of AR(1) in the mgcv package is sub-optimal, as it only allows for one $\rho$ value per model, effectively treating the entire dataset as one big time series. However, the degree of autocorrelation actually varies considerably across the tokens in our dataset. Applying a single $\rho$ correction will therefore under-correct for some tokens but over-correct for others, and could thus create an unwanted effect on the distribution of the residuals. We discuss this further in the supplementary material.
A standard method for evaluating the effects of different predictors in regression modeling is to compare nested models in which individual variables are progressively added or removed. However, this approach assumes that the predictors are relatively independent of one another, so that the contribution of each predictor to overall variance can be identified. If highly correlated predictors are entered into a single model, it becomes impossible to separate the effects of one from another and hence to assess their relative contributions. It is self-evident that the segmental makeup of a word cannot be independent of that word and that, in our models, the core predictor word is almost completely identified by the segment-related controls and vice versa. Thus, in a model that included both word and the controls, it would be impossible to evaluate the effects of the different predictors. The same applies to word and sense, since the latter is completely nested under the former. Consequently, the standard approach to model comparison was not available to us as a means of testing our predictions, since we were specifically aiming to tease apart the effects of different variables.
Instead of comparing nested models that would have been linguistically uninterpretable, we adopted the following two-fold modeling strategy. Firstly, in order to evaluate whether a given predictor was relevant for understanding F0 contours, we made use of Akaike’s Information Criterion (AIC, Akaike,, 1998). AIC estimates the relative quality of alternative models for a given set of data, by balancing goodness of fit against the number of parameters for each model, penalizing excessively complex models that might overfit the data. The method can be used for non-nested models, and is therefore well suited to our data. The best model is the one with the lowest value of AIC. [^15] For model comparison using AIC, the evidence ratio is calculated as $e^{\frac{\Delta AIC}{2}}$. A decrease of 10 AIC units, for example, means that the model with the smaller AIC is 148.4 times more likely than the alternative model to generate the observed data. Secondly, we investigated the adequacy of our predictors by means of cross-validation. That is, we held out a small portion of our dataset as testing data, fitted our models to the remaining data, and then assessed model accuracy on the testing data. We repeated this process for 100 different data splits selected by stratified random sampling such that every word type was represented in both the training and test data, in similar proportions. In other words, the test sets consisted entirely of novel tokens but no novel types, simulating the situation for human language use with previous experience. Using cross-validation, we could directly assess the precision with which our models predicted novel, previously unseen, data, and establish whether inclusion of a predictor improved prediction accuracy. [^16] The cross-validation approach we used is sometimes called repeated training/test split or Monte Carlo cross-validation. It is one of the most common cross-validation methods in machine learning (Zhang,, 1993; Kuhn,, 2013).
@@
    Word type will be a stronger predictor of tonal realization than all the previously established word-form related predictors combined.
@@
    Information about a word’s meaning in context will improve prediction of its tonal realization, compared with prediction based on word type alone.
@@
Prediction 1 is based on the hypothesis that the unique pitch contour of each spoken Mandarin word token is determined in part by the meaning of that token. If this is correct, then variations in tonal realization arise not only from the mechanical constraints on articulation previously described in the literature, but also as a result of form-meaning connections in the lexicon. We therefore investigated whether using word as a predictor would lead to a more precise model of tonal realization, as compared to a model using the set of segment-related predictors (e.g., vowel height) that have previously been identified as relevant. Since word incorporates not only the segmental form of a word type but also the associated semantics, our expectation was that it would be a superior predictor compared to all the previously identified segment-related variables considered jointly, when we controlled for contextual effects such as token duration and adjacent tones.
@@
As a first step, we fitted a baseline GAM to logarithmically transformed F0, including all the speaker-related and context-related control variables described in Sections 2.3.2 and 2.3.3, in interaction with time. We then fitted six further models, each with one segment-related control variable added to the baseline, allowing us to investigate the articulatory effects described in the literature. To compare these effects jointly against the effect of word, we fitted two additional models: one with all six segment-related control variables included in addition to the baseline, and the other with only word as an additional predictor. Both word and the segment-related controls (vowel1, vowel2, onset1, onset2, syllable1, syllable2) were modeled as factor smooths in interaction with time; in other words, all these variables have both random intercepts and random wiggly effects on the shapes of the predicted contours.
If Prediction 1 is correct, and we find an effect of word type, this will provide some support for the hypothesis that meaning contributes to the realization of tone. However, it will not enable us to draw any firm conclusion. This is because word type subsumes information about form as well as meaning. Even if word performs better than the combination of all previously identified segmental predictors, we cannot be certain that this difference is due to meaning; it is possible that some additional aspect of word form influences tonal realization but has not yet been identified as a predictor. However, at the word level, it is impossible to tease these effects apart. Since all our word types have the same RF tonal pattern, if we were to enter every detail of a token’s segmental makeup into the model, we would effectively be specifying its word type. Prediction 2 addresses this issue by separating variation in meaning from variation in form.
Prediction 2 is based on the hypothesis that variation in tonal realization is partly determined by token-level variation in meaning, i.e., variation in meaning within word type. To address this prediction, we compared the model with a factor smooth for word as the sole predictor added to the baseline against a model in which a factor smooth for sense was the sole predictor added to the baseline, both in interaction with time. Recall that, because the sense labels in our data are nested under the orthographic form, sense includes all the information in word, plus additional information about the meaning of any given token. If semantics is indeed at issue, replacing word by sense should therefore improve model fit even further.
@@
In the dataset for the word-type analysis, the frequency distribution of sense is skewed towards the right, with about half of the senses having no more than 13 tokens. To make sure that all senses included in the models had sufficient tokens for statistical evaluation, we therefore used only a subset of the data for the models used to investigate the effect of sense. Since the median number of tokens per sense in the dataset was 13.5, we excluded senses with fewer than 14 tokens. This left us with a dataset of 3,458 tokens representing 65 senses across a smaller set of 48 word types. We used this smaller dataset for models evaluating sense as a predictor. [^17] For completeness, we note that 35 words in the smaller dataset have only one sense. Of these 35 words, six have only one sense listed in the Chinese Wordnet, and seven are not included in the vocabulary of the Chinese WordNet. For the rest, either the tagger only identified one sense, or only one sense of the word had more than 13 tokens in our dataset. The statistical analysis proceeded in the same way as described above for word type, except that we added the additional model with a factor smooth for sense as the only predictor in addition to the baseline.
@@ 2.5 The baseline GAM
We fitted the baseline GAM to log-transformed F0 (pitch), using the following model specification: [^18] All model formulae and summaries are also provided in the supplementary material.
@@
s(duration) + ti(time, duration) +
@@
The first line of this model requests a main effect for gender, in order to account for male voices being lower on average than female voices. We used treatment (i.e. dummy) coding, with female as the reference level. In addition, the model requests a separate smooth for each gender. The upper left-hand panel of Figure 3 plots the predicted contours for speakers identified as female (red) and those identified as male (blue). Similar to the pattern observed in read speech (cf. Figure 1), the realization of the RF tonal pattern in spontaneous speech is characterized by a shallow fall, followed by a long rise, and finally a much larger fall. Speakers identified as male show reduced pitch excursion compared to those identified as female, presumably due to male voices having a more compressed pitch range. The second line of the model requests by-speaker nonlinear random effects, using factor smooths. These factor smooths specify, for each speaker, the specific way in which that particular speaker modulates the general F0 contour associated with their gender. [^19] The syntax bs=’fs’ on the second line of the baseline model specification has a similar effect to the by argument on the first line; both terms request a smooth for each level of a single factor variable (in this case, speaker and gender respectively). The by argument is generally preferred when the factor has few levels, and the levels are of interest, as in the case of gender. For computational efficiency, the fs argument is preferred when there are many levels and these levels are of less direct interest, as in the case of speaker. When specifying a by-smooth, a separate term requesting a main effect for the intercept needs to be specified. In contrast, a factor smooth incorporates adjustments to the intercept, thus effectively calibrating the individual smooths for their relative position with respect to the general intercept. See Baayen et al., (2022) for detailed discussion.
@@
The next four lines in the model formula deal with the four numerical context-related controls, namely duration, utterance_position, bigram_previous and bigram_following. For each of these variables, the model requests a main effect smooth in combination with a tensor product smooth for the interaction of the given variable with time (using an interaction-specific tensor product smooth specified with ti). The upper mid panel of Figure 3 plots the modulating effect of token duration on the base contour of speakers identified as female. Shorter duration, represented by darker shades of red, reduces the amplitude of the wave. The effect of position in the utterance is depicted in the upper right-hand panel of Figure 3. The tonal shape is clearly most different when the word occurs towards the end of an utterance, in which case we observe an earlier peak. This might be due to the fact that we coded words in singleton utterances as occurring at the end of the utterance. The left and mid panels in the lower row of Figure 3 present the effects of the bigram probabilities given the preceding and following word respectively; higher bigram probabilities are represented by lighter shades of red. When bigram_previous is high, meaning that the word is more expected given the preceding word, the F0 excursion is reduced. This effect parallels the finding of Hsieh, (2013) that F0 excursion in Taiwan Mandarin is diminished in conditions of high semantic predictability. The effect of bigram_following in our model is much smaller than the other contextual effects but appears to go in the opposite direction to bigram_previous, with higher values associated with sharper peaks in F0.
@@
Figure 3: Partial effects in the baseline GAM. The upper left-hand panel shows the predicted base contours for speakers identified as female (red) and speakers identified as male (blue). The next four panels show, for female speakers, how the base contour is modulated by duration, utterance position, previous bigram probability, and following bigram probability, respectively. The final panel presents, again for female speakers, the effect of tonal coarticulation with the tone of the preceding word, when the following word has a high-level tone.
@@ 2.6.1 Evaluation of predictors
The left-hand panel of Figure 4 presents the relative improvement in model fit, as compared to the baseline model, for the six models with an additional factor smooth for a single segment-related control, the model with factor smooths for all segment-related controls (henceforth the omnibus-segment model), and the model with only a factor smooth for word as additional predictor. Improvement in model fit is gauged by the magnitude of any decrease in AIC. As can be seen, each individual segment-related control improves model fit, a result that dovetails well with the previous studies summarized in Section 1.
In the omnibus-segment model, the inclusion of all six segment-related controls jointly provides a substantially better fit than any single predictor and leads to a fall in AIC of 4,938 AIC units compared to the baseline model. However, this is a very poor model from a statistical point of view because its key predictors are correlated with one another. In a GAM, the concurvity score of a predictor is a number bounded between zero and one that measures the degree to which the effect of a given independent variable can be predicted by one or more of the other independent variables in the model. [^20] Concurvity is the nonlinear equivalent of collinearity. If a predictor’s concurvity is low, this predictor has its own explanatory value; however, if the concurvity is high, the predictor is strongly confounded with other predictors. The blue dots in the right-hand panel of Figure 4 represent the concurvity scores of the segment-related controls in the omnibus-segment model. It can be seen that the concurvity scores of all predictors are high, indicating that the effects of the segment-related controls are confounded with one another, rendering interpretation of the individual effects difficult if not impossible. [^21] As explained in Section 2.4, this high level of concurvity, which would be further increased by the addition of word, is the reason why we cannot create a meaningful model that includes both word and the segment-related controls. Despite the fact that the omnibus-segment model is linguistically uninterpretable, we have presented it here purely for comparison with the model with only word as additional predictor.
@@
The concurvity score of the predictor word in the word-type model is represented by the red dot in the right-hand panel of Figure 4. It can be seen that word has low concurvity with the other predictors in this model (i.e., the baseline controls), so that the interpretation of the effect of individual word types on the F0 contours is straightforward. Furthermore, adding only word to the baseline model results in an even better fit than the omnibus-segment model, with a fall of 6,795 AIC units compared to the baseline. Although the segment-related controls address all the word-internal properties previously found to influence tonal realization, the contribution of just word by itself to the model fit is much stronger. The difference of 1,857 AIC units between the omnibus segment model and the word model means that the probability of the word model giving a better fit approaches infinity. It is clear that the association between word type and tonal realization cannot be reduced to the segment-related constraints on articulation previously described in the literature. The actual pitch contour is richer than what can be predicted from all phonetic features that have been found to be relevant. Note, however, that we do not claim that the segmental predictors are irrelevant, as our analyses clearly demonstrate that they all improve on the baseline model. We also do not claim that once word is included, the segmental predictors are irrelevant. What we do claim is that word by itself outperforms all segmental predictors jointly, and hence has added predictive value over and above the segmental predictors.
@@
Figure 4: The left-hand panel shows model fit improvement gauged by decrease in AIC units when a predictor (or set of predictors) is added to the baseline model for the word-type analysis. The right-hand panel shows the concurvity score of individual predictors in two models using the full dataset of 3,778 tokens: the omnibus-segment model (blue) with factor smooths for all segment-related control variables added to the baseline, and the word model (red) with only a factor smooth for word added to the baseline.
@@
The predicted pitch contours of a sample of 15 word types are presented in Figure 5. To better visualize how the word-specific tonal modulations differ from one another, the partial effect predicted for each word has been added to the general contour for speakers identified as female (cf. Figure 3). In general, the fall-rise-fall pattern can be observed for all these words, but the details of tonal excursions differ significantly from word to word. For example, while the initial falling part is very prominent for words like (c) {CJK*}UTF8bsmi決定 jue2ding4 ‘decision’ and (f) {CJK*}UTF8bsmi全部, quan2bu4 ‘all’, it is rather muted for (m) {CJK*}UTF8bsmi一半 yi2ban4 ‘half’ and (d) {CJK*}UTF8bsmi年紀 nian2ji4 ‘age’. In addition, in terms of the degree of undulation, some words have more reduced tonal range, such as (a) {CJK*}UTF8bsmi不是 bu2shi4 ‘not’ as compared to (j) {CJK*}UTF8bsmi文化 wen2hua4 ‘culture’, which has an extensive F0 excursion.
@@
Figure 5: Examples of the pitch contours predicted by the general smooth for time for female speakers, combined with the partial effects of the factor smooth for word. These partial effects do not include the general intercept, nor the differences in pitch between female and male speakers. As they represent the pure effect of word on the pitch contour, irrespective of other predictors, the curves are centered around the y-axis (indicated by a horizontal dotted line). The vertical dotted lines in the panels indicate the average (word-specific) syllable boundary.
@@
2.6.2 Cross-validation
Recall from Section 2.4 that the goal of the cross-validation analysis is to investigate whether the fitted model can make precise predictions for held-out data, i.e. tokens that were withheld from the model during model fitting (training). If our hypothesis is correct, a GAM that has access to word should provide superior prediction accuracy on held-out data compared to a GAM that has access only to the segment-related controls. We therefore evaluated prediction accuracy under cross-validation. We held out 10% of the current data as test data, and used the remaining 90% as training data. Every word type was represented in both the training data and the test data, with approximately the same distribution in each set. We used stratified random sampling to produce 100 different splits with these properties.